# When is an Embedder More Promising than Another?

### v1

@@ B.3 Proof of Proposition 2 and Corollary 1
$\max\limits_{x\in\mathsf{X}}\sum\limits_{y\in\mathsf{Y}}\left|\displaystyle\int_{\mathcal{A}_{y}}P_{Z|X}(\mathsf{d}z|x)-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)\right|\leqslant 2\varepsilon|\mathsf{X}|,$
for all partitions of $\mathsf{Z}$. This concludes the proof of the Corollary 1. ∎
@@ B.4 Example of comparisons of statistical experiments
Let $U|x$ and $Z|x$ be independently normally distributed as $\mathcal{N}(x,\sigma^{2})$ and $\mathcal{N}(x,\epsilon^{2}\sigma^{2})$ , respectively, with $10<\epsilon<$ 1.
  • Case of $\sigma^{2}=\sigma^{2}_{0}$ known. Here $U\succcurlyeq_{S}Z$ since $Z+\nu|x$ has the same distribution as $U|x$ when $\nu\sim\mathcal{N}(0,(1-\epsilon^{2})\sigma_{0}^{2})$. That is $Z$ is strictly more informative than $U$. However, $U$ is strictly more informative than $Z$.
  • Case of $x=0$ known. One can observe that $U\approx Z$ since the variables $Z/\epsilon$ have the same distribution as the $U$, and the variables $\epsilon U$ have the same distribution as the $Z$.
  • Case of $x$ and $\sigma$ unknown. Surprisingly, in this case $U$ and $Z$ are not comparable.
@@ B.5 Sufficiency and Inference Procedures with Embedding Models [13, 63]
An embedding model $P_{U|Y}\in\mathcal{K}(\mathsf{U}|\mathsf{Y})$ is deemed to be sufficient for another one $P_{Z|Y}\in\mathcal{K}(\mathsf{Z}|\mathsf{Y})$ if and only if, for any bounded loss function $\ell$ where $\|\ell\|_{\infty}\leqslant 1$, and for any inference procedure $\rho_{Z}:\mathsf{Z}\rightarrow\mathsf{Y}$, there exists a inference procedure (possibly randomized) $\rho_{U}:\mathsf{U}\rightarrow\mathcal{P}(\mathsf{Y})$ such that the resulting statistical risks satisfy
$\mathcal{R}_{y}(P_{U|Y},\rho_{U},\ell)\leqslant\mathcal{R}_{y}(P_{Z|Y},\rho_{Z},\ell),\quad\textrm{ for all $y\in\mathsf{Y}$}.$ (21)
Here we denote by $\mathcal{R}_{y}(P_{U|Y},\rho_{Y},\ell)$ and $\mathcal{R}_{y}(P_{Z|Y},\rho_{Z},\ell)$ the statistical risks for the corresponding inference frameworks, respectively.
@@ B.6 Deficiency and Expected Risk [63]
Let $\varepsilon>0$ be fixed. Then, $\delta(P_{U|Y}\rightarrow P_{Z|Y})<\varepsilon$ if and only if, for any bounded loss function $\ell$ where $\|\ell\|_{\infty}\leqslant 1$, and for any inference procedure $\rho_{Z}$ using the embedding model $P_{Z|Y}$, there exists a inference procedure (possibly randomized) $\rho_{U}$ based the embedding model $P_{U|Y}$ such that the risks satisfy $\mathcal{R}_{y}(P_{U|Y},\rho_{U},\ell)-\varepsilon\leqslant\mathcal{R}_{y}(P_{Z|Y},\rho_{Z},\ell),\quad\textrm{ for all $y\in\mathsf{Y}$}.$
@@ C.1 Models and Datasets statistics
Salesforce/SFR-Embedding-Mistral 4096 32768 0.59
jspringer/echo-mistral-7b-instruct-lasttoken 4096 32768 0.58
infgrad/stella-base-en-v2 768 512 0.57
intfloat/e5-large-v2 1024 512 0.57
GritLM/GritLM-7B 4096 32768 0.56
llmrails/ember-v1 1024 512 0.56
thenlper/gte-large 1024 512 0.56
WhereIsAI/UAE-Large-V1 1024 512 0.55
thenlper/gte-base 768 512 0.55
jamesgpt1/sf_model_e5 1024 512 0.55
avsolatorio/GIST-Embedding-v0 768 512 0.55
sentence-transformers/gtr-t5-large 768 512 0.55
sentence-transformers/gtr-t5-xl 768 512 0.55
BAAI/bge-base-en-v1.5 768 512 0.54
sentence-transformers/sentence-t5-large 768 512 0.54
google/gemma-7b-it N/A N/A 0.54
TaylorAI/gte-tiny 384 512 0.53
sentence-transformers/gtr-t5-base 768 512 0.53
NousResearch/Llama-2-7b-hf N/A N/A 0.53
sentence-transformers/sentence-t5-xl 768 512 0.53
google/gemma-2b-it N/A N/A 0.52
intfloat/e5-small 384 512 0.52
SmartComponents/bge-micro-v2 N/A N/A 0.51
sentence-transformers/all-distilroberta-v1 N/A N/A 0.51
intfloat/multilingual-e5-small 384 512 0.51
sentence-transformers/msmarco-bert-co-condensor 768 512 0.51
princeton-nlp/sup-simcse-bert-base-uncased 768 512 0.50
sentence-transformers/all-MiniLM-L6-v2 384 512 0.50
sentence-transformers/all-mpnet-base-v2 768 514 0.49
izhx/udever-bloom-560m 1024 2048 0.49
sentence-transformers/LaBSE 768 512 0.47
sentence-transformers/average_word_embeddings_komninos 300 N/A 0.42
sentence-transformers/average_word_embeddings_glove.6B.300d 300 N/A 0.41
sentence-transformers/allenai-specter 768 512 0.38
@@ C.3.1 Full MTEB Benchmark Results
Salesforce/SFR-Embedding-Mistral 67.56 78.33 51.67 60.64 59.00 85.05
jspringer/echo-mistral-7b-instruct-lasttoken 64.68 77.43 46.32 58.14 55.52 82.56
infgrad/stella-base-en-v2 62.61 75.28 44.90 58.78 50.10 83.02
intfloat/e5-large-v2 62.25 75.24 44.49 56.61 50.56 82.05
GritLM/GritLM-7B 66.76 79.46 50.61 60.49 57.41 83.35
llmrails/ember-v1 63.54 75.99 45.58 60.04 51.92 83.34
thenlper/gte-large 63.13 73.33 46.84 59.13 52.22 83.35
WhereIsAI/UAE-Large-V1 64.64 75.58 46.73 59.88 54.66 84.54
thenlper/gte-base 62.39 73.01 46.20 58.61 51.14 82.30
jamesgpt1/sf_model_e5 63.34 73.96 46.61 59.86 51.80 83.85
avsolatorio/GIST-Embedding-v0 63.71 76.03 46.21 59.37 52.31 83.51
sentence-transformers/gtr-t5-large 58.28 67.14 41.60 55.36 47.42 78.19
sentence-transformers/gtr-t5-xl 58.42 67.11 41.51 55.96 47.96 77.80
BAAI/bge-base-en-v1.5 63.55 75.53 45.77 58.86 53.25 82.40
sentence-transformers/sentence-t5-large 57.06 72.31 41.65 54.00 36.71 81.83
google/gemma-7b-it N/A N/A N/A N/A N/A N/A
TaylorAI/gte-tiny 58.69 70.35 42.09 55.77 44.92 80.46
sentence-transformers/gtr-t5-base 56.19 65.25 38.63 54.23 44.67 77.07
NousResearch/Llama-2-7b-hf N/A N/A N/A N/A N/A N/A
sentence-transformers/sentence-t5-xl 57.87 72.84 42.34 54.71 38.47 81.66
google/gemma-2b-it N/A N/A N/A N/A N/A N/A
intfloat/e5-small 58.89 71.67 39.51 54.45 46.01 80.87
SmartComponents/bge-micro-v2 N/A N/A N/A N/A N/A N/A
sentence-transformers/all-distilroberta-v1 N/A N/A N/A N/A N/A N/A
intfloat/multilingual-e5-small 57.87 70.74 37.08 53.87 46.64 79.10
sentence-transformers/msmarco-bert-co-condensor 52.35 64.71 37.64 51.84 32.96 76.47
princeton-nlp/sup-simcse-bert-base-uncased 48.87 67.32 33.43 47.54 21.82 79.12
sentence-transformers/all-MiniLM-L6-v2 56.26 63.05 42.35 58.04 41.95 78.90
sentence-transformers/all-mpnet-base-v2 57.78 65.07 43.69 59.36 43.81 80.28
izhx/udever-bloom-560m 55.81 68.04 36.89 52.60 41.19 79.93
sentence-transformers/LaBSE 45.21 62.71 29.55 48.42 18.99 70.80
sentence-transformers/average_word_embeddings_komninos 42.06 57.65 26.57 44.75 21.22 62.46
sentence-transformers/average_word_embeddings_glove.6B.300d 41.96 57.29 27.73 43.29 21.62 61.85
sentence-transformers/allenai-specter 40.28 52.37 34.06 48.10 15.88 61.02
@@ E.1 Estimation method
Embedding dimension normalization.
@@
Median instead of mean.
@@ E.4 Impact of the number of models
E.5 Computational ressources

### v2

@@ B.3 Proof of Proposition 2 and Corollary 1
$\sup\limits_{\{A_{y}\}}\,\mathbb{E}\sum\limits_{y\in\mathsf{Y}}\left|\displaystyle\int_{\mathcal{A}_{y}}P_{V|X}(\mathsf{d}v|X)-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|X)\right|\leqslant 2|\mathsf{Y}|\varepsilon.$
This concludes the proof of the Corollary 1. ∎
@@ B.4 Example of comparisons of statistical experiments
Let $U|x$ and $V|x$ be independently normally distributed as $\mathcal{N}(x,\sigma^{2})$ and $\mathcal{N}(x,\epsilon^{2}\sigma^{2})$ , respectively, with $10<\epsilon<$ 1.
  • Case of $\sigma^{2}=\sigma^{2}_{0}$ known. Here $U\succcurlyeq_{S}V$ since $V+\nu|x$ has the same distribution as $U|x$ when $\nu\sim\mathcal{N}(0,(1-\epsilon^{2})\sigma_{0}^{2})$. That is $V$ is strictly more informative than $U$. However, $U$ is strictly more informative than $V$.
  • Case of $x=0$ known. One can observe that $U\approx V$ since the variables $V/\epsilon$ have the same distribution as the $U$, and the variables $\epsilon U$ have the same distribution as the $V$.
  • Case of $x$ and $\sigma$ unknown. Surprisingly, in this case $U$ and $V$ are not comparable.
@@ B.5 Sufficiency and Inference Procedures with Embedding Models [13, 63]
An embedding model $P_{U|Y}\in\mathcal{K}(\mathsf{U}|\mathsf{Y})$ is deemed to be sufficient for another one $P_{V|Y}\in\mathcal{K}(\mathsf{V}|\mathsf{Y})$ if and only if, for any bounded loss function $\ell$ where $\|\ell\|_{\infty}\leqslant 1$, and for any inference procedure $\rho_{V}:\mathsf{V}\rightarrow\mathsf{Y}$, there exists a inference procedure (possibly randomized) $\rho_{U}:\mathsf{U}\rightarrow\mathcal{P}(\mathsf{Y})$ such that the resulting statistical risks satisfy
$\mathcal{R}_{y}(P_{U|Y},\rho_{U},\ell)\leqslant\mathcal{R}_{y}(P_{V|Y},\rho_{V},\ell),\quad\textrm{ for all $y\in\mathsf{Y}$}.$ (25)
Here we denote by $\mathcal{R}_{y}(P_{U|Y},\rho_{Y},\ell)$ and $\mathcal{R}_{y}(P_{V|Y},\rho_{V},\ell)$ the statistical risks for the corresponding inference frameworks, respectively.
@@ B.6 Deficiency and Expected Risk [63]
Let $\varepsilon>0$ be fixed. Then, $\delta(P_{U|Y}\rightarrow P_{V|Y})<\varepsilon$ if and only if, for any bounded loss function $\ell$ where $\|\ell\|_{\infty}\leqslant 1$, and for any inference procedure $\rho_{V}$ using the embedding model $P_{V|Y}$, there exists a inference procedure (possibly randomized) $\rho_{U}$ based the embedding model $P_{U|Y}$ such that the risks satisfy $\mathcal{R}_{y}(P_{U|Y},\rho_{U},\ell)-\varepsilon\leqslant\mathcal{R}_{y}(P_{V|Y},\rho_{V},\ell),\quad\textrm{ for all $y\in\mathsf{Y}$}.$
@@ C.1 Models and Datasets statistics
SFR-Embedding-Mistral 4096 32768 0.59
echo-mistral-7b-instruct-lasttoken 4096 32768 0.58
stella-base-en-v2 768 512 0.57
e5-large-v2 1024 512 0.57
GritLM-7B 4096 32768 0.56
ember-v1 1024 512 0.56
gte-large 1024 512 0.56
UAE-Large-V1 1024 512 0.55
gte-base 768 512 0.55
sf_model_e5 1024 512 0.55
GIST-Embedding-v0 768 512 0.55
gtr-t5-large 768 512 0.55
gtr-t5-xl 768 512 0.55
bge-base-en-v1.5 768 512 0.54
sentence-t5-large 768 512 0.54
gemma-7b-it N/A N/A 0.54
gte-tiny 384 512 0.53
gtr-t5-base 768 512 0.53
Llama-2-7b-hf N/A N/A 0.53
sentence-t5-xl 768 512 0.53
gemma-2b-it N/A N/A 0.52
e5-small 384 512 0.52
bge-micro-v2 384 512 0.51
all-distilroberta-v1 N/A N/A 0.51
multilingual-e5-small 384 512 0.51
msmarco-bert-co-condensor 768 512 0.51
sup-simcse-bert-base-uncased 768 512 0.50
all-MiniLM-L6-v2 384 512 0.50
all-mpnet-base-v2 768 514 0.49
udever-bloom-560m 1024 2048 0.49
LaBSE 768 512 0.47
average_word_embeddings_komninos 300 N/A 0.42
average_word_embeddings_glove.6B.300d 300 N/A 0.41
allenai-specter 768 512 0.38
@@ C.3.1 Full MTEB Benchmark Results
SFR-Embedding-Mistral 67.56 78.33 51.67 60.64 59.00 85.05
echo-mistral-7b-instruct-lasttoken 64.68 77.43 46.32 58.14 55.52 82.56
stella-base-en-v2 62.61 75.28 44.90 58.78 50.10 83.02
e5-large-v2 62.25 75.24 44.49 56.61 50.56 82.05
GritLM-7B 66.76 79.46 50.61 60.49 57.41 83.35
ember-v1 63.54 75.99 45.58 60.04 51.92 83.34
gte-large 63.13 73.33 46.84 59.13 52.22 83.35
UAE-Large-V1 64.64 75.58 46.73 59.88 54.66 84.54
gte-base 62.39 73.01 46.20 58.61 51.14 82.30
sf_model_e5 63.34 73.96 46.61 59.86 51.80 83.85
GIST-Embedding-v0 63.71 76.03 46.21 59.37 52.31 83.51
gtr-t5-large 58.28 67.14 41.60 55.36 47.42 78.19
gtr-t5-xl 58.42 67.11 41.51 55.96 47.96 77.80
bge-base-en-v1.5 63.55 75.53 45.77 58.86 53.25 82.40
sentence-t5-large 57.06 72.31 41.65 54.00 36.71 81.83
gemma-7b-it N/A N/A N/A N/A N/A N/A
gte-tiny 58.69 70.35 42.09 55.77 44.92 80.46
gtr-t5-base 56.19 65.25 38.63 54.23 44.67 77.07
Llama-2-7b-hf N/A N/A N/A N/A N/A N/A
sentence-t5-xl 57.87 72.84 42.34 54.71 38.47 81.66
gemma-2b-it N/A N/A N/A N/A N/A N/A
e5-small 58.89 71.67 39.51 54.45 46.01 80.87
bge-micro-v2 56.57 68.04 39.18 54.29 42.56 78.65
all-distilroberta-v1 N/A N/A N/A N/A N/A N/A
multilingual-e5-small 57.87 70.74 37.08 53.87 46.64 79.10
msmarco-bert-co-condensor 52.35 64.71 37.64 51.84 32.96 76.47
sup-simcse-bert-base-uncased 48.87 67.32 33.43 47.54 21.82 79.12
all-MiniLM-L6-v2 56.26 63.05 42.35 58.04 41.95 78.90
all-mpnet-base-v2 57.78 65.07 43.69 59.36 43.81 80.28
udever-bloom-560m 55.81 68.04 36.89 52.60 41.19 79.93
LaBSE 45.21 62.71 29.55 48.42 18.99 70.80
average_word_embeddings_komninos 42.06 57.65 26.57 44.75 21.22 62.46
average_word_embeddings_glove.6B.300d 41.96 57.29 27.73 43.29 21.62 61.85
allenai-specter 40.28 52.37 34.06 48.10 15.88 61.02
@@ E.1 Estimation method
E.1.1 KNIFE Estimator
We recall the definition of the information sufficiency:
$\mathcal{I}_{S}(U\rightarrow Z)\triangleq\underbrace{\inf_{f\in\mathcal{F}_{\Theta}(\mathsf{Z})}\mathbb{E}\left[-\log f(Z)\right]}_{\text{Uncertainty of }Z}-\underbrace{\mathbb{E}\left[\inf_{M\in\mathcal{K}_{\Theta}(\mathsf{Z}|\mathsf{U})}\mathbb{E}\left[-\log M(Z|U)|U\right]\right]}_{\text{Uncertainty when simulating }Z\text{ from }U\text{ with }M}.$
We denote $C$ the number of modes chosen for the Gaussian mixture distributions and $\operatorname{GM}_{\bm{\mu,\Sigma,w}}$ the Gaussian mixture distribution with $C$ components, parametrized by $\bm{\mu,\Sigma,w}$, with $\mathbf{1}^{T}\bm{w}=1$, such that
$\operatorname{GM}_{\bm{\mu,\Sigma,w}}(z)=\sum_{c=1}^{C}\bm{w}_{c}\mathcal{N}(z|\bm{\mu}_{c},\bm{\Sigma}c)$, where $\mathbf{w}_{c}$ is the weight of the $c$-th component, and $\mathcal{N}(z|\bm{\mu}_{c},\bm{\Sigma}_{c})$ is the density of a multivariate Gaussian distribution with mean $\bm{\mu}_{c}$ and covariance $\bm{\Sigma}_{c}$ the c-th mean and covariance matrix.
To estimate the information sufficiency between the two embedders $U$ and $Z$, we follow the procedure described in KNIFE [82].
$\mathcal{F}_{\Theta}$ is hence the class of multivariate Gaussian mixtures with $C$ components, it is parametrized by $\bm{\mu,\Sigma,w}$. These learnable parameters are optimized to maximize the log-likelihood of $Z$ [^8] In practice since this operation does not depend on $U$, we store the weights of this distribution, and use it for all pairs $\mathcal{I}_{S}(.\rightarrow Z)$.
The class of Markov kernels $\mathcal{K}_{\Theta}(\mathsf{Z}|\mathsf{U})$ is also composed of multivariate Gaussian mixtures whose parameters are estimated using a small feedforward neural network. Such that for each $u\in\mathsf{U}$, the parameters of the Gaussian mixture are $\bm{\mu}(u),\bm{\Sigma}(u),\bm{w}(u)$.
In practice we considered the covariance matrix to be diagonal to avoid the number of parameters of the Gaussian mixtures to grow too large with the dimension of the embeddings $d$. The number of parameters to be estimated for the Gaussian mixtures are hence: $C\times(2d+1)$, with the number of parameters of the feedforward networks for the Markov kernels.
Procedure 1 Estimation of $\mathcal{I}_{S}(U\rightarrow Z)$, $\operatorname{GM}_{\mathbf{\mu},\mathbf{\Sigma},\mathbf{w}}$ denotes the Gaussian Mixture model with means $\mathbf{\mu}$, covariances $\mathbf{\Sigma}$ and weights $\mathbf{w}$.
0:  Pairs of corresponding embeddings $(z_{i},u_{i})_{N}$
0:  Information sufficiency $\mathcal{I}_{S}(U\rightarrow Z)$
  $\mathbf{\mu_{Z}},\mathbf{\Sigma_{Z}},\mathbf{w_{Z}}\leftarrow\arg\min_{\mathbf{\mu},\mathbf{\Sigma},\mathbf{w}}-\sum_{i=1}^{N}\log\operatorname{GM}_{\mathbf{\mu},\mathbf{\Sigma},\mathbf{w}}(z_{i})$
  $\mathbf{\mu_{Z|U}},\mathbf{\Sigma_{Z|U}},\mathbf{w_{Z|U}}\leftarrow\arg\min_{\mathbf{\mu},\mathbf{\Sigma},\mathbf{w}}-\sum_{i=1}^{N}\ log\operatorname{GM}_{\mathbf{\mu}(u_{i}),\mathbf{\Sigma}(u_{i}),\mathbf{w}(u_{i})}(z_{i})$
  Uncertainty of $Z$: $H(Z)\leftarrow\frac{1}{N}\sum_{i=1}^{N}\log\operatorname{GM}_{\mathbf{\mu_{Z}},\mathbf{\Sigma_{Z}},\mathbf{w_{Z}}}(z_{i})$
  Uncertainty of $Z$ given $U$: $H(Z|U)\leftarrow\frac{1}{N}\sum_{i=1}^{N}\log\operatorname{GM}_{\mathbf{\mu_{Z|U}}(u_{i}),\mathbf{\Sigma_{Z|U}}(u_{i}),\mathbf{w_{Z|U}}(u_{i})}(z_{i})$
  Return $\mathcal{I}_{S}(U\rightarrow Z)\leftarrow H(Z)-H(Z|U)$
Both the parameters of the marginal distribution of $Z$ and the parameters of the conditional distribution of $Z$ given $U$ are estimated through likelihood maximization. The uncertainty of $Z$ is estimated by the negative log-likelihood of the data under the marginal distribution of $Z$. The uncertainty of $Z$ given $U$ is estimated by the negative log-likelihood of the data under the conditional distribution of $Z$ given $U$.
E.1.2 Embedding dimension normalization.
@@ E.1.1 KNIFE Estimator
E.1.3 Median instead of mean.
@@ E.4 Impact of the number of models
E.5 Comparison with other metrics
We evaluate other ways to measure the informativeness (Tab. 9 and Tab. 10) of the embedders. We considered the size of the model, their embeding dimensions and the simple reconstruction error when we train a cross-encoder to go from one embedding to another. Overall we found that our information sufficiency score significantly outperformed these naive metrics in both modalities.
In particular, training a cross-encoder with the $l_{2}$ reconstruction error proves to correlate with the ability to enable downstream performance, however, this correlation remains weaker compared to using the information sufficiency.
Table 9: Comparison with Baselines: Size of the Embedder, Dimension of the embedding output ($d$) and the $\ell_{2}$ reconstruction error of the embeddings for NLP datasets.
Size d $\bar{I}_{\mathbf{S}}$$\bar{\ell}_{2}$
$\varrho_{p}$ $\varrho_{s}$ $\tau$ $\varrho_{p}$ $\varrho_{s}$ $\tau$ $\varrho_{p}$ $\varrho_{s}$ $\tau$ $\varrho_{p}$ $\varrho_{s}$ $\tau$
Classification (12 datasets) 0.46 0.42 0.32 0.52 0.66 0.55 0.92 0.88 0.73 -0.79 -0.85 -0.66
Retrieval (15 datasets) 0.40 0.39 0.29 0.46 0.63 0.52 0.89 0.89 0.70 -0.71 -0.84 -0.65
Clustering (11 datasets) 0.45 0.38 0.26 0.54 0.67 0.55 0.86 0.85 0.67 -0.80 -0.84 -0.66
STS (10 datasets) 0.27 0.35 0.25 0.34 0.66 0.52 0.92 0.82 0.62 -0.70 -0.83 -0.64
Reranking (4 datasets) 0.33 0.33 0.26 0.41 0.61 0.50 0.84 0.79 0.64 -0.71 -0.78 -0.59
Average (56 datasets) 0.41 0.41 0.31 0.48 0.62 0.50 0.94 0.90 0.74 -0.77 -0.84 -0.65
Additional Classif (8 datasets) 0.41 0.62 0.47 0.43 0.64 0.55 0.89 0.84 0.66 -0.65 -0.72 -0.55
Table 10: Comparison with Baselines: Size of the Embedder, Dimension of the embedding output ($d$) and the $\ell_{2}$ reconstruction error of the embeddings for Molecular Modeling datasets.
Size d $\bar{I}_{\mathbf{S}}$$\bar{\ell}_{2}$
$\varrho_{p}$ $\varrho_{s}$ $\tau$ $\varrho_{p}$ $\varrho_{s}$ $\tau$ $\varrho_{p}$ $\varrho_{s}$ $\tau$ $\varrho_{p}$ $\varrho_{s}$ $\tau$
Absorption (8 datasets) - -0.21 -0.16 - -0.43 -0.29 - 0.89 0.70 - -0.89 -0.70
Distribution (3 datasets) - -0.07 -0.03 - -0.46 -0.31 - 0.89 0.70 - -0.86 -0.66
Metabolism (8 datasets) - 0.06 0.03 - -0.46 -0.34 - 0.94 0.79 - -0.90 -0.71
Excretion (3 datasets) - -0.17 -0.11 - -0.24 -0.18 - 0.77 0.60 - -0.77 -0.56
Toxicity (9 datasets) - 0.09 0.06 - -0.49 -0.35 - 0.92 0.75 - -0.86 -0.67
ADMET (31 datasets) - -0.01 0.01 - -0.47 -0.32 - 0.94 0.80 - -0.90 -0.72
E.6 Computational ressources